# The Relative Entropy of Expectation and Price

### v7

@@ I Introduction
Advocates of deep hedging replace risk neutralisation with risk optimisation as the guiding principle for pricing, as data-driven models see the complex correlations among economic variables as opportunity rather than hindrance. Risk optimisation is similarly used here to derive the price measure from the expectation measure, though the objectives differ. Model risk is quantified as the residual risk after funding and hedging, which requires for its definition a precise understanding of the relationship between expectation and price. Taking this idea to its logical conclusion, the convex risk metric is implemented as margin held against model risk. Optimisation of the strategic return net of all settlements then naturally embeds risk optimisation.
@@ III.1 Calibration
which is solved with the unit optimal portfolio portfolio:
@@
which is solved with the unit optimal portfolio portfolio:
@@ III.2 Hedging and pricing
These conditions express a non-linear relationship between the initial and final derivative prices. By definition, the derivative security with notional $\lambda$ has final price $\lambda P$. The corresponding initial price $p$, hedge portfolio $\delta$ and implied funding spread $s$ per unit notional are determined from the hedge and price conditions:
@@ V The Black-Scholes model
$-\frac{1}{2}\alpha q^{2}\nu\left((\omega-\delta)-\left(\frac{\mu-r}{\alpha q\nu}-\frac{\partial p}{\partial q}\right)\right)^{\!2}+O[dt]$
@@
$-\frac{1}{2}\alpha q^{2}\nu\left((\omega-\delta)-\left(\frac{\mu-r}{\alpha q\nu}-\frac{\partial p}{\partial q}-\frac{\partial p}{\partial\nu}\frac{\rho\epsilon}{q}\right)\right)^{\!2}$
@@ VII Discrete models
$\mathbb{E}[X]=\frac{1}{n}\operatorname{tr}[X]$ (143)
When optimising the strategy with this target function, care must be taken to manage the potential complications of non-commutativity. For example, the perturbation $\partial(X^{2})=(\partial X)X+X(\partial X)$ cannot be simplified further when $X$ and $\partial X$ do not commute. Fortunately, the trace of this expression can be simplified, thanks to cyclicity of the trace, and in general the perturbation has the handy simplification:
The calibration, hedge and price conditions are then derived from entropic risk optimisation as:
$0$ $=\operatorname{tr}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (146)
$0$ $=\operatorname{tr}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]$
$p$ $=-\frac{1}{\alpha}\log\!\left[\frac{\operatorname{tr}[\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]}{\operatorname{tr}[\exp[-\alpha\omega\cdot dq]]}\right]$
These conditions are explored in the studies of figures˜11, 12, 13 and 14, assuming an underlying market that contains a funding security with initial price $b=1$ and final price $B=b+db$ satisfying $\operatorname{tr}[db]=0$ and a hedging security with initial price $q$ and final price $Q=q+dq$ satisfying $\operatorname{tr}[Q]=0$. The model is developed for the final derivative price expressed as a function P=f[QkB]P=f[Q-kB] of the spread QkBQ-kB at strike $k$.
VII.1 Classical study
The classical studies of figures˜11, 12 and 13 assume $B=1$ and $Q$ is a diagonal matrix with eigenvalues $Q_{i}$ sampled from a standard normal variable, centralised to have zero sample mean. The matrix P=f[QkB]P=f[Q-kB] is then diagonal with eigenvalues $P_{i}:=f[Q_{i}-k]$.
@@ VII.1 Classical study
$0=\sum_{i=1}^{n}(Q_{i}-q)\exp[-\alpha\omega Q_{i}]$ (147)
is solved for the optimal portfolio $\omega$, which generates the eigenstate probabilities:
$\gamma_{i}:=\frac{\displaystyle\exp[-\alpha\omega Q_{i}]}{\displaystyle\sum\nolimits_{j=1}^{n}\exp[-\alpha\omega Q_{j}]}$ (148)
@@
VII.2 Quantum study
The quantum studies of figure˜14 assume $q=0$ and $B$ and $Q$ are matrices with $n=2$ eigenvalues:
@@ VII.2 Quantum study
The calibration condition is trivially solved with $\omega=0$, and the hedge and price conditions become:
@@ VIII Entropic XVA
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$\lambda=0$ and $c=0.5$ $\lambda=1$ and $c=0.5$ $\lambda=2$ and $c=0.5$
@@ VIII.2 Entropic XVA
To understand the XVA contributions, consider the simplified model with predictable funding:
@@
where $r_{1}$ and $r_{2}$ are the funding rates of the underlying securities and $r$ is the funding rate of the margin account. With only one security in each underlying market, the self-funding conditions derive $\delta_{1}=c_{1}/q_{1}$ and $\delta_{2}=c_{2}/q_{2}$ and the threshold prices become:
$p_{1}={}$ $\frac{1}{1+r_{1}\,dt}\mathbb{E}_{1}^{\phi_{1}}[\alpha_{1}]\!\left[P_{-m_{1}(1+r\,dt)}^{m_{2}(1+r\,dt)}\right]$ (159)
$-\frac{m_{1}(r_{1}-r)\,dt}{1+r_{1}\,dt}$
$p_{2}={}$ $\frac{1}{1+r_{2}\,dt}\mathbb{E}_{2}^{\phi_{2}}[-\alpha_{2}]\!\left[P_{-m_{1}(1+r\,dt)}^{m_{2}(1+r\,dt)}\right]$
$+\frac{m_{2}(r_{2}-r)\,dt}{1+r_{2}\,dt}$
where $P=p+dp$ is the final derivative price and the margins are $m_{1}:=c_{1}-p$ and $m_{2}:=c_{2}+p$ for the two counterparties. The first term is the discounted entropy-adjusted mean in the price measure of the final derivative price, capped and floored by the respective margins to generate the credit and debit valuation adjustments. The second term then adjusts the initial price by the funding valuation adjustment.

### v8

@@ I Introduction
Advocates of deep hedging replace risk neutralisation with risk optimisation as the guiding principle for pricing, as data-driven models see the complex correlations among economic variables as opportunity rather than hindrance. Risk optimisation is similarly used here to derive the price measure from the expectation measure, though the objectives differ. Model risk is quantified as the residual risk after funding and hedging, which requires for its definition a precise understanding of the relationship between expectation and price. Taking this idea to its logical conclusion, the convex risk metric is implemented as margin held against model risk. Optimisation of the strategic return net of funding and hedging then adjusts price to account for the risk management benefits of margin.
@@ III.1 Calibration
which is solved with the unit optimal portfolio:
@@
which is solved with the unit optimal portfolio:
@@ III.3 Limit order book
These conditions express a non-linear relationship between the initial and final derivative prices. If the derivative security with notional $\lambda$ has final price $\lambda P$, then the corresponding initial price $p$, hedge portfolio $\delta$ and implied funding spread $s$, quoted per unit notional of the derivative security, are determined from the hedge and price conditions:
@@ V The Black-Scholes model
$-\frac{1}{2}\alpha q^{2}\nu\left(\left(\omega-\frac{\mu-r}{\alpha q\nu}\right)-\left(\delta-\frac{\partial p}{\partial q}\right)\right)^{\!2}+O[dt]$
@@
$-\frac{1}{2}\alpha q^{2}\nu\left(\left(\omega-\frac{\mu-r}{\alpha q\nu}\right)-\left(\delta-\frac{\partial p}{\partial q}-\frac{\partial p}{\partial\nu}\frac{\rho\epsilon}{q}\right)\right)^{\!2}$
@@ VII Discrete models
$\mathbb{E}[X]:=\frac{1}{n}\operatorname{tr}[X]$ (143)
When optimising the strategy with this target function, care must be taken to manage the potential complications of non-commutativity. For example, the perturbation $\partial(X^{2})=(\partial X)X+X(\partial X)$ cannot be simplified further when $X$ and $\partial X$ do not commute. Fortunately, the trace of this expression can be simplified, thanks to cyclicity, and in general the perturbation has the handy simplification:
Using this result, the portfolios are optimised at the stationary points of the entropy-adjusted mean:
$0$ $=\frac{\partial}{\partial\omega}(\mathbb{E}[\alpha][\omega\cdot dq]-\omega\cdot qr\,dt)$ (146)
$=\frac{\displaystyle\operatorname{tr}[dq\,\exp[-\alpha\omega\cdot dq]]}{\displaystyle\operatorname{tr}[\exp[-\alpha\omega\cdot dq]]}-qr\,dt$
$0$ $=\frac{\partial}{\partial\delta}(\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]-(p-\delta\cdot q)(r+\alpha s)\,dt)$
$=-\frac{\displaystyle\operatorname{tr}[dq\,\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]}{\displaystyle\operatorname{tr}[\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]}+q(r+\alpha s)\,dt$
which generates the calibration and hedge conditions. The price condition for the initial derivative price is then obtained by equating the entropy-adjusted means. With the unit optimal portfolio defined by $\phi:=\alpha\omega$, these three conditions are expressed as:
$0$ $=\operatorname{tr}[(dq-qr\,dt)\exp[-\phi\cdot dq]]$ (147)
$0$ $=\operatorname{tr}[(dq-q(r+\alpha s)\,dt)\exp[-\phi\cdot dq-\alpha(P-\delta\cdot dq)]]$
$p$ $=-\frac{1}{\alpha}\log\!\left[\frac{\operatorname{tr}[\exp[-\phi\cdot dq-\alpha(P-\delta\cdot dq)]]}{\operatorname{tr}[\exp[-\phi\cdot dq]]}\right]$
The calibration condition is solved together with the self-funding condition $\phi\cdot q=0$ for the implied funding rate $r$ and the unit optimal portfolio $\phi$, and the hedge and price conditions are solved together with the self-funding condition $\delta\cdot q=p$ for the implied funding spread $s$, the hedge portfolio $\delta$ and the initial derivative price $p$.
VII.1 Mid pricing
There are, in general, no explicit solutions to the hedge and price conditions, though they are efficiently solved by the Newton-Raphson scheme. Mid pricing is derived in the limit $\alpha=0$, and for this special case the conditions are explicitly solved to derive the hedge portfolio and the initial derivative price. The solution depends on the small-$\alpha$ expansion of the conditions, which is complicated by the non-commutativity of the underlying returns. Fortunately, the expansion can be generated with the aid of Poincaré’s formula for the differential of the matrix exponential.
Define the adjoint action $\operatorname{ad}_{X}[Y]:=XY-YX$ of the matrix $X$ on the matrix $Y$. Poincaré’s formula is derived in the form:
$1$-{} $\exp[-(X+\alpha Y)]\exp[X]$ (148)
$=\int_{s=0}^{1}\frac{\partial}{\partial s}(1-\exp[-s(X+\alpha Y)]\exp[sX])\,ds$
$=\alpha\int_{s=0}^{1}\exp[-s(X+\alpha Y)]Y\exp[sX]\,ds$
$=\alpha\int_{s=0}^{1}\exp[-sX]Y\exp[sX]\,ds+O[\alpha^{2}]$
$=\alpha\left(\int_{s=0}^{1}\exp[-s\operatorname{ad}_{X}]\,ds\right)\![Y]+O[\alpha^{2}]$
$=\alpha\frac{\displaystyle 1-\exp[-\operatorname{ad}_{X}]}{\displaystyle\operatorname{ad}_{X}}[Y]+O[\alpha^{2}]$
where the foundational relationship between Lie algebras and Lie groups:
$\exp[\operatorname{ad}_{X}][Y]=\exp[X]Y\exp[-X]$ (149)
is applied at the fourth step. Combining this result with the elementary property $\operatorname{tr}[\operatorname{ad}_{X}[Y]Z]=-\operatorname{tr}[Y\operatorname{ad}_{X}[Z]]$ of the adjoint action derives the expansion:
$\operatorname{tr}[Z\exp$ $[-(X+\alpha Y)]]=\operatorname{tr}[Z\exp[-X]]$ (150)
$-\alpha\operatorname{tr}\!\left[\frac{\displaystyle\exp[\operatorname{ad}_{X}]-1}{\displaystyle\operatorname{ad}_{X}}[Z]Y\exp[-X]\right]+O[\alpha^{2}]$
This result is used to solve the hedge and price conditions in the mid-pricing case. Define the price measure:
$\mathbb{E}^{\phi}[X]:=\frac{\displaystyle\operatorname{tr}[X\exp[-\phi\cdot dq]]}{\displaystyle\operatorname{tr}[\exp[-\phi\cdot dq]]}$ (151)
Poincaré’s formula derives the small-$\alpha$ expansions:
$\frac{\operatorname{tr}[\exp[-\phi\cdot dq-\alpha(P-\delta\cdot dq)]]}{\operatorname{tr}[\exp[-\phi\cdot dq]]}=$ (152)
$\qquad\qquad 1-\alpha\mathbb{E}^{\phi}[P-\delta\cdot dq]+O[\alpha^{2}]$
$\frac{\operatorname{tr}[dq\,\exp[-\phi\cdot dq-\alpha(P-\delta\cdot dq)]]}{\operatorname{tr}[\exp[-\phi\cdot dq]]}=$
$\qquad\qquad qr\,dt-\alpha\mathbb{E}^{\phi}[d\hat{q}\,(P-\delta\cdot dq)]+O[\alpha^{2}]$
adjusting to account for non-commutativity in the underlying returns with the definition:
$d\hat{q}:=\frac{\displaystyle\exp[\operatorname{ad}_{\phi\cdot dq}]-1}{\displaystyle\operatorname{ad}_{\phi\cdot dq}}[dq]$ (153)
The expansions generate the following solutions for the hedge and price conditions:
$p$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[P]+O[\alpha]$ (154)
$s\,dt$ $=\frac{\displaystyle p-\mathbb{V}^{\phi}[d\hat{q},dq]^{-1}\mathbb{V}^{\phi}[d\hat{q},P]\cdot q}{\displaystyle\mathbb{V}^{\phi}[d\hat{q},dq]^{-1}q\cdot q}+O[\alpha]$
$\delta$ $=\mathbb{V}^{\phi}[d\hat{q},dq]^{-1}(\mathbb{V}^{\phi}[d\hat{q},P]+qs\,dt)+O[\alpha]$
Quantum solutions are similar to the corresponding classical solutions, but include a correction to the hedge portfolio when there is non-commutativity among the components of the underlying return vector.
VII.2 Option pricing
Consider a market that contains a funding security whose final price is represented by the positive-definite matrix $B$ and an underlying security whose final price is represented by the self-adjoint matrix $Q$. For the spread portfolio with strike $k$ whose final price is represented by the self-adjoint matrix QkBQ-kB, let $u$ be an eigenstate and let $\zeta$ be the corresponding eigenvalue satisfying:
$(Q-kB)u=\zeta u$ (155)
The sensitivity of the eigenvalue to the strike is negative:
$\frac{d\zeta}{dk}=-\frac{\displaystyle u^{\dagger}Bu}{\displaystyle u^{\dagger}u}$ (156)
The eigenstate may also be sensitive to the strike, and this creates higher-order convexity in the relationship between strike and eigenvalue. If the matrices $Q$ and $B$ commute then the eigenstates of the matrix QkBQ-kB are fixed as the strike varies and the corresponding eigenvalues are linear in the strike. If the matrices $Q$ and $B$ do not commute then the eigenstates of the matrix QkBQ-kB are not fixed as the strike varies and the corresponding eigenvalues are non-linear in the strike.
Convexity between strike and eigenvalue can be utilised in option pricing. Assume for convenience that, for each $i=1,\ldots,n$, the $i$th eigenvalue $\zeta_{i}$ is strictly decreasing as a function of the strike and is zero at the unique strike $x_{i}$. If the calibration condition is solved with unit optimal portfolio $\phi=0$ and implied funding rate $r=0$, so that funding adjustments can be ignored, then the mid price for the option is:
$\frac{1}{n}\operatorname{tr}[(Q-kB)^{+}]=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}[k]^{+}$ (157)
$\qquad=\sum_{i=1}^{n}(x_{i}-k)^{+}w_{i}+\int_{x=-\infty}^{\infty}(x-k)^{+}w[x]\,dx$
where the second expression is derived from the first expression using integration by parts with:
$w_{i}$ $:=-\frac{1}{n}\zeta_{i}^{\prime}[x_{i}]$ (158)
$w[x]$ $:=\frac{1}{n}\sum_{i=1}^{n}\zeta_{i}^{\prime\prime}[x](1_{x<x_{i}})$
The option price is expressed as an integral of the payoff $(x-k)^{+}$ with discrete weight $w_{i}$ at the point $x_{i}$ and continuous density $w[x]$ at the point $x$. In the classical model with commuting matrices, the eigenvalues satisfy $\zeta_{i}^{\prime\prime}[x]=0$ and the implied distribution is discrete. In the quantum model with non-commuting matrices, the eigenvalues satisfy $\zeta_{i}^{\prime\prime}[x]\neq 0$ and the implied distribution has discrete and continuous components.
This ability of the model to extract an implied distribution with continuous support from discrete variables is the source of its efficacy. Heisenberg uncertainty can be calibrated to realistic market volatility smiles with as few as two eigenstates, demonstrating convergence in the quantum model that is exponentially faster than classical Monte-Carlo estimation. The eigenvalue convexity this exploits is precisely the feature that enables violation of the classical Bell’s inequality, and is a uniquely quantum property.
The quantum contribution to the model is captured in the unitary matrix $U$ that rotates from the eigenbasis of the matrix $Q$ to the eigenbasis of the matrix $B$. Starting with diagonal matrices for $Q$ and $B$, the option price is then expressed as:
$\frac{1}{n}\operatorname{tr}[(UQU^{-1}-kB)^{+}]$ (159)
No rotation between the eigenbases is generated when $U$ is given by the identity:
$U_{jk}=1_{j=k}$ (160)
For the simplest case $n=2$ this derives the option price:
$\frac{1}{2}(Q_{1}-kB_{1})^{+}+\frac{1}{2}(Q_{2}-kB_{2})^{+}$ (161)
Maximum rotation between the eigenbases is generated when $U$ is given by the discrete Fourier transform:
$U_{jk}=\frac{1}{\sqrt{n}}\exp\!\left[2\pi i\frac{jk}{n}\right]$ (162)
For the simplest case $n=2$ this derives the option price:
$\frac{1}{2}\!\left((\bar{Q}-k\bar{B})+\sqrt{\hat{Q}^{2}+k^{2}\hat{B}^{2}}\right)^{\!+}$ (163)
$+\frac{1}{2}\!\left((\bar{Q}-k\bar{B})-\sqrt{\hat{Q}^{2}+k^{2}\hat{B}^{2}}\right)^{\!+}$
where:
$\bar{B}$ $:=\frac{1}{2}(B_{1}+B_{2})$ $\qquad\hat{B}$ $:=\frac{1}{2}(B_{1}-B_{2})$ (164)
$\bar{Q}$ $:=\frac{1}{2}(Q_{1}+Q_{2})$ $\qquad\hat{Q}$ $:=\frac{1}{2}(Q_{1}-Q_{2})$
The additional convexity in the relationship between strike and eigenvalue creates an implied distribution with continuous support.
@@ VII.2 Option pricing
Generalisations of this model are explored in the studies of figures˜10, 11, 12, 13 and 14. The model is developed for the final derivative price expressed as a function P=f[QkB]P=f[Q-kB] of the spread portfolio QkBQ-kB at strike $k$, for the call option with $f[x]=x^{+}$ and the digital option with $f[x]=1_{x>0}$.
VII.3 Classical study
The classical studies of figures˜10, 11, 12 and 13 assume $b=1$ and $B=1$, and $Q$ is a diagonal matrix with eigenvalues $Q_{i}$ sampled from a standard normal variable, centralised to have zero sample mean. The matrix P=f[QkB]P=f[Q-kB] is then diagonal with eigenvalues $P_{i}:=f[Q_{i}-k]$.
@@ VII.3 Classical study
$0=\sum_{i=1}^{n}(Q_{i}-q)\exp[-\phi Q_{i}]$ (165)
is solved for the unit optimal portfolio $\phi$, which generates the eigenstate probabilities:
$\gamma_{i}:=\frac{\displaystyle\exp[-\phi Q_{i}]}{\displaystyle\sum\nolimits_{j=1}^{n}\exp[-\phi Q_{j}]}$ (166)
@@
VII.4 Quantum study
The quantum studies of figure˜14 assume $b=1$ and $q=0$, and $B$ and $Q$ are matrices with $n=2$ eigenvalues:
@@ VII.4 Quantum study
The calibration condition is trivially solved with $\phi=0$, and the hedge and price conditions become:
@@ VIII.2 Entropic XVA
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$\lambda=0$ and $c=0.5$ $\lambda=1$ and $c=0.5$ $\lambda=2$ and $c=0.5$
@@
To understand the XVA contributions, consider the simplified model with predictable funding and no hedging:
@@
where $r_{1}$ and $r_{2}$ are the funding rates of the funding securities and $r$ is the funding rate of the margin account. With only one security in each underlying market, the self-funding conditions derive $\delta_{1}=c_{1}/q_{1}$ and $\delta_{2}=c_{2}/q_{2}$ as well as $\phi_{1}=\phi_{2}=0$, and the threshold prices become:
$p_{1}={}$ $\frac{1}{1+r_{1}\,dt}(\mathbb{E}_{1}[\alpha_{1}][\hat{P}]-m_{1}(r_{1}-r)\,dt)$ (177)
$={}$ $\frac{1}{1+r_{1}\,dt}(\mathbb{E}_{1}[\hat{P}]-m_{1}(r_{1}-r)\,dt)$
$-\frac{\alpha_{1}}{2}\frac{1}{1+r_{1}\,dt}\mathbb{V}_{1}[\hat{P}]+O[\alpha_{1}^{2}]$
$p_{2}={}$ $\frac{1}{1+r_{2}\,dt}(\mathbb{E}_{2}[-\alpha_{2}][\hat{P}]+m_{2}(r_{2}-r)\,dt)$
$={}$ $\frac{1}{1+r_{2}\,dt}(\mathbb{E}_{2}[\hat{P}]+m_{2}(r_{2}-r)\,dt)$
$+\frac{\alpha_{2}}{2}\frac{1}{1+r_{2}\,dt}\mathbb{V}_{2}[\hat{P}]+O[\alpha_{2}^{2}]$
where $m_{1}:=c_{1}-p$ and $m_{2}:=c_{2}+p$ are the initial margins for the two counterparties. The final payoff is capped and floored by the default events:
$\hat{P}:=P_{-m_{1}(1+r\,dt)}^{m_{2}(1+r\,dt)}$ (178)
which generates the credit and debit valuation adjustments. For each counterparty, the threshold price is the discounted expectation of the capped and floored payoff, adjusted by the funding valuation adjustment and including a reserve against the unhedged risk of the payoff.